OpenStudy (anonymous):

1/2 log3 (x)-2 log3 (y)+4 log3 (z)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I hope you can read it well

OpenStudy (anonymous):

= log3x^(1/2) - log3 y^2 + log3 (z)^4 = log3 [(x^(1/2) z^4) / y^2]

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

is that - logs to base 3 (call that L) with 0.5L(x) - 2L(y) + 4L(z)? not sure, running out of time, but you might find these rules useful: AlogB = log(B^A) logA + log B = log(AB) log A - log B = log (A/B)

OpenStudy (anonymous):

Thank you for your help too. Those rules I needed thanks